# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v8

@@ 2.2 Tsukishima: A Representative Manufacturing Area
The population in Tokyo City has been growing since the end of the 19th century, and industrialization due to WWI spurred its growth. Tokyo City’s population share increased from 2.5% in 1889 to approximately 4% by 1920.161616In 1920, there were more than 10,000 municipalities in Japan, suggesting how huge Tokyo was. There was a difference of about 1 million between Tokyo and Osaka, the second-largest city by population. Online Appendix Figure A.1a summarizes the trends in the number of people in the top five cities between 1889 and 1920. Throughout the 1910s, the percentage of industrial and commercial workers increased in Tokyo City, building the city as a gigantic commercial and industrial place (Tanimoto 2002, p. 8). Column 1 of Panel A in Table 1 shows that nearly half of the male workers in Tokyo worked in the industrial sector as of 1920.
@@
Notes: The name in each lattice indicates the administrative ward in Tokyo. The gray ward indicates the Kyobashi Ward and an island in the Kyobashi Ward indicates Tsukishima. The border surrounded by Asakusa, Honjyo, Nihonbashi, and Fukagawa wards includes the Sumida River. The borders of each ward are based on the administrative district in 1920. Geographical coordinate system is based on JGD2000/(B, L).
@@
Notes: Blocks in green and light green indicate the shipyards. Blocks in red and pink indicate the factories in the machinery sector. Blocks in blue and light blue indicate the factories in the other manufacturing sectors. Blocks in brown are the warehouses. Small circles colored in black and white show the workplaces in the machinery sector and the other manufacturing sectors, respectively.
@@
The subject of this paper is Tsukishima, an industrial area located southeast of Kyobashi Ward in Tokyo (Figure 2). [^17] The base of the island was reclaimed during the Edo and Meiji periods from a delta at the mouth of the Sumida River. Online Appendix A.3 summarizes the development history of Tsukishima in detail. Column (2) of Panel A in Table 1 shows that Kyobashi had an industrial composition similar to that of Tokyo City in 1920. Column (3) confirms that Tsukishima is a representative industrial area in the city; approximately 68% of male workers were employed in the manufacturing sector. In the same year, there were 214 factories on this island, and approximately 80% of them belonged to the machinery and related sectors (Department of Health, Ministry of the Interior 1923a, pp 385–389). [^18] According to the census of 1920, the number of workers in the metal and machinery equipment manufacturing industry ($27.1$%) was the highest compared to other industries in the entire city of Tokyo. The same tendency was observed in Kyobashi Ward ($35.5$%). Although similar statistics are not available for Tsukishima, this supports the evidence that Kyobashi was not an area with a unique industrial pattern (Statistics Bureau of the Cabinet 1929c, pp. 85–86; 108–109)). Figure 3 shows the spatial distribution of factories in Tsukishima as of June 30, 1918. Tsukishima had many machine factories and shipyards, and the workers’ residences were clustered there. [^19] For example, a representative shipbuilding plant was located on Ishikawajima, an island in the north of Tsukishima (Figure 3). Ships were strongly influenced by the import substitution that was underway at the beginning of the 20th century, and domestic production of ships was progressing (Section 2.1). The production of a hydraulic 80-ton hardened crane by Ishikawajima Shipyard in 1917 is an example of the domestic output of a production good that had previously relied on imports (Nakamura and Odaka 1989, pp. 10; 32). The focus of this study is the consumption behavior of these worker households in the industrial sector.
@@
Panel A: This panel summarizes the industrial structures based on the occupations of male workers measured in the 1920 population census. Each share is calculated as the number of males worked as regular workers in each sector divided by the total number of male workers (in percentage). All the figures do not include the number of unemployed (mugyō sha) males.
Panel B: The average household size is the number of people divided by the number of households. The share of married males is the number of married males living with/without spouses divided by the total number of males (in percentage). The average age of males is calculated using the tables for population by age group reported in the census. Since the tables are open-ended, the class for over 60 years old is rounded as $64.5$ following the range of the former category (i.e., 50–59). All the figures in this panel do not include a small number of quasi-households (jyun setai), which include the person, such as the students in dormitories and the patients in hospitals.
@@
To look more closely at the characteristics of working-class households in Tsukishima, Panel B in Table 1 summarizes the population and occupational statistics from the census. First, the average household size in Tsukishima is $4.3$, similar but slightly smaller than the values in Tokyo City ($4.6$) and Kyobashi Ward ($4.7$). The percentage of the married male population is $40$%, which is in the similar range of the figures of Tokyo ($38$%) and Kyobashi ($36$%). The average age of men in Tsukishima is $25.1$, slightly lower than the figures for Tokyo and Kyobashi. While the sex ratio by age group in Tsukishima shows a similar trend to that of Tokyo and Kyobashi, the ratios for the 1414141419191919 and 2020202029292929 age bins are higher than those. Overall, these statistics reflect the fact that Tsukishima is a typical industrial area and many young male workers were employed there. The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory worker households. The percentage of married men in Tsukishima suggests that a certain proportion of the households in their early 20s might have been couples before they started raising children.
@@
Finally, I briefly overview the lives of the factory workers’ households in Tsukishima. Factories and workers’ residences were scattered throughout the island. [^20] Tsukishima is composed of Ishikawajima island at the northern end, the adjacent Tsukudajima and Shin-tsukudajima islands, and the rectangular Lots 1–3 in the south (Online Appendix Figure A.3). Specifically, factories and worker residences were scattered throughout Shin-tsukudajima and Lots 1 and 2, whereas Lot 3, a relatively new reclaimed site completed in 1913, was undeveloped. The factory workers had lived in similar tenements and used the commercial stores and restaurants in the central lot. [^21] Online Appendix A.4 summarizes the housing in Tsukishima in detail using a housing survey. Regarding the stores, there were 209 stores selling daily necessities, 109 selling clothing, 74 selling tools, and 120 selling food, drink, and entertainment. As for medical personnel, there were eight doctors, four dentists, eight obstetricians, and 14 drug stores. There were also ten bathhouses (Department of Health, Ministry of the Interior 1923a, pp. 41–47). There were two postal offices, 13 pawnshops, and five moneylenders, and those were scattered throughout the island (Online Appendix Figure A.2). The main routes to the mainland were the wooden Aioi Bridge (Aioi bashi) and three ferries (watashi bune). [^22] See Online Appendix Figure A.3 for these routes. The average daily total number of people using the ferries in 1918 was reported to be $53,318$, accounting for more than twice as many people as the Tsukishima population. [^23] See Department of Health, Ministry of the Interior (1923a, p. 30). The number of people and households of Tsukishima at the end of December 1918 is reported to be $24,399$ and $5,562$, respectively (National Police Agency 1920a, pp. 50–51). These descriptions illustrate Tsukishima as a busy machine industry area during the industrialization.
@@ 2.3 Risk-Coping Institutions
Working-class households had been exposed to several risks that led to losses of income at that time. Generally, social insurance, such as health, accident, and unemployment insurance, can help mitigate the idiosyncratic income shocks on the workers. [^24] Gertler and Gruber (2002) shows the impacts of illness on consumption smoothing behavior in early 1990s Indonesia and argues the importance of public support for the illness. Kantor and Fishback (1996) show that the installation of workers’ compensation reduced the precautionary savings in the interwar U.S.. Gruber (1997) provides evidence that unemployment insurance smooths individual consumption as the case of the U.S. from 1968–1987. However, comprehensive social insurance systems did not exist in the early 1920s in Japan. The former health insurance was enacted in 1927, and unemployment insurance had not been established throughout the prewar period. In addition, the Factory Act enacted in 1916 did also not specify compensation for the sickness of workers. Therefore, the workers had basically no public support for their illness, injuries, and joblessness. Instead, they might have used private financial institutions to deal with the idiosyncratic risks.
@@
Several financial institutions were indeed available for working-class households. A representative savings institution for households was the postal savings (yūbin chokin). In Tokyo City, roughly $1.4$ million people had postal savings accounts, covering approximately $67$% of the entire citizens in 1920 (Online Appendix Table A.2). The average savings amount per capita in Kyobashi Ward was $44.5$ yen (Tokyo City Office 1921, pp. 892–893), which is less than one month’s earnings of the skilled-factory workers (Section 3.2). Savings bank (chochiku ginkō) is another famous savings institution among the working class in the 1910s (Ito and Saito 2019, pp. 79–81). The number of people with savings accounts was roughly $1.4$ million, accounting for approximately $58$% of total people in Tokyo (Online Appendix Table A.2). However, $84.8$% of all the depositors were the commerce and miscellaneous workers, and the manufacturing sector shared only $11.5$% in Kyobashi Ward (Online Appendix Table A.3). This reflects the fact that savings bank was an institution designed for lower-income workers than factory workers (Tokyo Institute for Municipal Research 1925b, p. 81). This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices (Section 2.2). Although another savings institution was the mutual loan association (mujin), it had been used among small business owners and merchants for running their businesses (Online Appendix A.2). [^25] Generally, the ordinary bank in prewar Japan was the institution for providing large loans to businesses (Teranishi 2011). There was no bank in Tsukishima, as workers did not borrow money from banks to make ends meet. The official report states that “a commercial bank is not necessary in a laboring area such as this Tsukishima, and even for large entrepreneurs such as factory owners, there is no particular need to establish a bank for their business transactions, and transactions with banks in other areas are sufficient” (Department of Health, Ministry of the Interior 1923a, p. 48).
@@
There also be a few lending institutions. Credit purchases (kakegai) in retail shops were the most popular institution. Although the systematic statistics on credit purchases are unavailable, the THBS households often used credit purchases in various retail shops for rice, fish, vegetables, firewood, and charcoal. [^26] Cooperative societies were also available for worker households to buy daily commodities, albeit they were not lending institutions. There were 26 cooperative societies in Tokyo prefecture in 1924, of which 23 were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). However, while the two largest societies had roughly $4,000$ members, the others were generally small societies with less than 100 members. This figure accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand. Most of the THBS households had not used the cooperatives. I will describe this institution in detail in Section 5. The pawnshop (shichiya) was another lending institution. The accessibilities were higher than the other lending institutions: [^27] The number of pawnshops in Tokyo were $1,334$ in 1918 and $1,261$ in 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. there were $13$ pawnshops in Tsukishima, more than twice the large number of moneylenders (Section 2.2). However, the core users of the pawnshops were the low-income working-class households rather than the factory workers (Tokyo City Social Affairs Bureau 1921, p. 9). According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% of all users. In contrast, those for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and $15.6$%, respectively (Tokyo Institute for Municipal Research 1926, p. 25). The inexpensive article means that pawnshops were used to deal with very short-run necessities only. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which accounts for $12$% of the monthly disposable income of the THBS households (Panel A of Table 3). In this light, the pawnshop is a small lending institution for low-income working-class households who might have faced credit constraints. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the main articles were inexpensive clothes (Shibuya et al. 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates. [^28] This means that the average redemption rate was substantially high: it was approximately $94$% in 1920 among $88$ pawnshops in Kyobashi ($401,222$ cases/$427,265$ cases) (Tokyo City Office 1922b, pp. 888–889).
@@
Other lending institutions are money lenders (kinsen kashitsuke gyō) and credit unions (shinyō kumiai). Generally, moneylenders were used by business owners with very high-interest rates (Shibuya 2000, pp. 184; 248). Similarly, the coverage of the credit unions was substantially low, say a few percentage points of the male workers in the manufacturing sector (Online Appendix A.2). Informal gifts may be another type of risk-coping institution, although systematic statistics are not available. A household survey on the 185 factory-worker households in Tokyo suggests that the share of average monthly gifts in total income was approximately $3$% in November 1922 (Social Affairs Division 1925, pp. 58-59). While this means that the gifts comprised a few percentage points of the monthly income, it is unclear whether these informal income transfers had functioned as mutual aid.
@@
Finally, life insurance was also not designed to compensate for temporary income losses. Generally, life insurance companies provided death insurance, meaning that temporary income reductions were not covered by insurance. Moreover, they did not target the working-class as customers (Usami 1984, pp. 87–88). Although postal life insurance (kani seimei hoken) operated by the government started in 1916, it only provided whole life insurance and endowment insurance. Thus, it also never covered the health shocks nor layoffs (Usami 1984, pp. 102–107).  [^29] There were also several employment agencies that introduced jobs to low-income job-seekers in Tokyo. However, a large part of them was unscrupulous and not commonly used at that time (Machida 2016).
@@
To summarize, while there were several institutions, the available institutions for the factory workers were few in the practical sense. For savings institutions, the postal savings might have been the best available device for Tsukishima workers. Regarding lending institutions, credit purchase is the plausible device among factory workers. While pawnshops were accessible financial institutions, they might not provide sufficient money as the articles were usually inexpensive. The factory workers in Tsukishima might have combined these devices to deal with the idiosyncratic income shocks.

### v10

@@ 2.2 Tsukishima: A Representative Manufacturing Area
The population in Tokyo City has been growing since the end of the 19th century, and industrialization due to WWI spurred its growth. Tokyo City’s population share increased from $2.5$% in 1889 to approximately $4$% by 1920.141414In 1920, there were more than $10,000$ municipalities in Japan, suggesting how large Tokyo was. There was a difference of about 1 million between Tokyo and Osaka, the second-largest city by population. Online Appendix Figure A.1a summarizes the trends in the number of people in the top five cities between 1889 and 1920. Throughout the 1910s, the percentage of industrial and commercial workers increased in Tokyo City, building the city as a gigantic commercial and industrial place (Tanimoto 2002, p. 8). Column 1 of Panel A in Table 1 shows that nearly half of the male workers in Tokyo worked in the industrial sector as of 1920.
@@
Notes: The name in each lattice indicates the administrative ward in Tokyo. The gray ward indicates the Kyobashi Ward and an island in the Kyobashi Ward indicates Tsukishima. The border surrounded by Asakusa, Honjyo, Nihonbashi, and Fukagawa wards includes the Sumida River. The borders of each ward are based on the administrative districts in 1920. Geographical coordinate system is based on JGD2000/(B, L).
@@
Notes: Blocks in green and light green indicate the shipyards. Blocks in red and pink indicate the factories in the machinery sector. Blocks in blue and light blue indicate the factories in the other manufacturing sectors. Blocks in brown are the warehouses. Small circles colored in black and white show the workplaces in the machinery and other manufacturing sectors, respectively.
@@
The subject of this paper is Tsukishima, an industrial area located southeast of Kyobashi Ward in Tokyo (Figure 2). [^15] The base of the island was reclaimed during the Edo and Meiji periods from a delta at the mouth of the Sumida River. Online Appendix A.3 summarizes the development history of Tsukishima in detail. Column (2) of Panel A in Table 1 shows that Kyobashi had an industrial composition like that of Tokyo City in 1920. Column (3) confirms that Tsukishima is a representative industrial area in the city; approximately $68$% of male workers were employed in the manufacturing sector. In the same year, there were $214$ factories on this island, and approximately $80$% of them belonged to the machinery and related sectors (Department of Health, Ministry of the Interior 1923a, pp 385–389). [^16] According to the census of 1920, the number of workers in the metal and machinery equipment manufacturing industry ($27.1$%) was the highest compared to other industries in the entire city of Tokyo. The same tendency was observed in Kyobashi Ward ($35.5$%). Although similar statistics are not available for Tsukishima, this supports the evidence that Kyobashi was not an area with a unique industrial pattern (Statistics Bureau of the Cabinet 1929c, pp. 85–86; 108–109)). Figure 3 shows the spatial distribution of factories in Tsukishima as of June 30, 1918. Tsukishima had many machine factories and shipyards, and the workers’ residences were clustered there. [^17] For example, a representative shipbuilding plant was located on Ishikawajima, an island in the north of Tsukishima (Figure 3). Ships were strongly influenced by the import substitution that was underway at the beginning of the 20th century, and domestic production of ships progressed (Section 2.1). The production of a hydraulic 80-ton hardened crane by Ishikawajima Shipyard in 1917 is an example of the domestic output of a product that had previously relied on imports (Nakamura and Odaka 1989, pp. 10; 32). The focus of this study is the consumption behavior of these worker households in the industrial sector.
@@
Panel A: This panel summarizes the industrial structures based on the occupations of male workers measured in the 1920 population census. Each share is calculated as the number of males who worked as regular workers in each sector divided by the total number of male workers (in percentage). All the figures do not include the number of unemployed (mugyō sha) males.
Panel B: The average household size is the number of people divided by the number of households. The share of married males is the number of married males living with/without spouses divided by the total number of males (in percentage). The average age of males is calculated using the tables for population by age group reported in the census. As the tables are open-ended, the class for over 60 years old is rounded as $64.5$ following the range of the former category (i.e., 50–59). All the figures in this panel do not include a small number of quasi-households (jyun setai), which include the person, such as the students in dormitories and the patients in hospitals.
@@
Panel B in Table 1 summarizes the population and occupational statistics from the census to examine the characteristics of working-class households in Tsukishima. First, the average household size in Tsukishima is $4.3$—similar to but slightly smaller than the values in Tokyo City ($4.6$) and Kyobashi Ward ($4.7$). The percentage of the married male population is $40$%, which is similar to the range of the figures for Tokyo ($38$%) and Kyobashi ($36$%). The average age of men in Tsukishima is $25.1$, slightly lower than the figures for Tokyo and Kyobashi. While the sex ratio by age group in Tsukishima shows a similar trend to that of Tokyo and Kyobashi, the ratios for the 1414141419191919 and 2020202029292929 age bins are higher than those. Overall, these statistics reflect the fact that Tsukishima was a typical industrial area and many young male workers were employed there. The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory worker households. The percentage of married men in Tsukishima suggests that a certain proportion of the households in their early 20s might have been couples before they started raising children.
@@
Finally, I provide a brief overview of the lives of the factory workers’ households in Tsukishima. Factories and workers’ residences were scattered throughout the island. The factory workers lived in similar tenements and used the commercial stores and restaurants in the central lot. [^18] Tsukishima is composed of Ishikawajima island at the northern end, the adjacent Tsukudajima and Shin-tsukudajima islands, and the rectangular Lots 1–3 in the south (Online Appendix Figure A.3). Specifically, factories and worker residences were scattered throughout Shin-tsukudajima and Lots 1 and 2, whereas Lot 3, a relatively new reclaimed site completed in 1913, was undeveloped. Online Appendix A.4 summarizes the housing in Tsukishima in detail using a housing survey. There were $209$, $109$, $74$, and $120$ stores selling daily necessities, clothing, tools, and food, drink, and entertainment, respectively. Regarding medical personnel, there were eight doctors, four dentists, eight obstetricians, and $14$ drug stores. There were also ten bathhouses (Department of Health, Ministry of the Interior 1923a, pp. 41–47). There were two postal offices, $13$ pawnshops, and five moneylenders, which were scattered throughout the island (Online Appendix Figure A.2). The main routes to the mainland were the wooden Aioi Bridge (Aioi bashi) and three ferries (watashi bune). The average daily total number of people using the ferries in 1918 was reported to be $53,318$, accounting for more than twice as many people as the Tsukishima population. [^19] See Department of Health, Ministry of the Interior (1923a, p. 30). The number of people and households in Tsukishima at the end of December 1918 was reported to be $24,399$ and $5,562$, respectively (National Police Agency 1920a, pp. 50–51). Online Appendix Figure A.3 shows the routes of the ferries. These descriptions illustrate Tsukishima as a busy machine manufacturing area during industrialization.
@@ 2.3 Risk-coping Institutions
Working-class households were exposed to several risks that led to losses of income at that time. Generally, social insurance, such as health, accident, and unemployment insurance, can mitigate idiosyncratic income shocks on workers. [^20] Gertler and Gruber (2002) shows the impact of illness on consumption-smoothing behavior in the early 1990s in Indonesia and argues for the importance of public compensation for illness. Kantor and Fishback (1996) show that the installation of workers’ compensation reduced the precautionary savings in the interwar US. Gruber (1997) provides evidence that unemployment insurance smooths individual consumption as was the case in the US from 1968–1987. However, comprehensive social insurance systems did not exist in the early 1920s in Japan. The previous health insurance was enacted in 1927, and unemployment insurance was not established throughout the prewar period. Additionally, the Factory Act enacted in 1916 did not specify compensation for the sickness of workers. Therefore, they had no public support for their illness, injuries, and unemployment. Instead, they might have used private financial institutions to manage idiosyncratic risks.
@@
Several financial institutions were available for working-class households. A representative savings institution was the postal savings (yūbin chokin). In Tokyo City, approximately $1.4$ million people had postal savings accounts, covering approximately $67$% of the citizens in 1920 (Online Appendix Table A.2). The average savings amount per capita in Kyobashi Ward was $44.5$ yen (Tokyo City Office 1921, pp. 892–893), which was less than one month’s earnings for the skilled-factory workers (Section 3.2). Savings bank (chochiku ginkō) was another famous savings institution among the working class in the 1910s (Ito and Saito 2019, pp. 79–81). The number of people with savings accounts was approximately $1.4$ million, accounting for approximately $58$% of total people in Tokyo (Online Appendix Table A.2). However, $84.8$% of all the depositors were commercial and miscellaneous workers, and the manufacturing sector shared only $11.5$% in Kyobashi Ward (Online Appendix Table A.3). This reflects the fact that savings banks were institutions designed for lower-income workers rather than factory workers (Tokyo Institute for Municipal Research 1925b, p. 81). This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices (Section 2.2). Although another savings institution was the mutual loan association (mujin), it was used among small business owners and merchants for running their businesses (Online Appendix A.2). [^21] Ordinary banks in prewar Japan provided large loans to businesses (Teranishi 2011). There was no bank in Tsukishima, as workers did not borrow money from banks to make ends meet. The official report states that “a commercial bank is not necessary in a laboring area such as Tsukishima, and even for large entrepreneurs such as factory owners, there is no particular need to establish a bank for their business transactions; transactions with banks in other areas are sufficient” (Department of Health, Ministry of the Interior 1923a, p. 48).
@@
There were also few lending institutions. Credit purchases (kakegai) with retailers were the most popular institution. Although systematic statistics of credit purchases are unavailable, the THBS households often used credit purchases in various retailers for rice, fish, vegetables, firewood, and charcoal. [^22] Cooperative societies were also available for worker households to buy daily commodities, although they were not lending institutions. There were $26$ cooperative societies in Tokyo prefecture in 1924, of which 23 were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). However, while the two largest societies had approximately $4,000$ members, the others were generally small societies with less than 100 members. This figure accounts for only a few percentage of the workers in the city, given that the number of male workers was more than $768$ thousand. In fact, the THBS households rarely used the cooperatives. I will describe this institution in detail in Section 5. The pawnshop (shichiya) was another lending institution. The accessibility was higher than for the other lending institutions. [^23] The number of pawnshops in Tokyo was $1,334$ and $1,261$ in 1918 and 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. There were $13$ pawnshops in Tsukishima, more than twice the number of moneylenders (Section 2.2). However, the primary users of the pawnshops were the low-income working-class households rather than the factory workers (Tokyo City Social Affairs Bureau 1921, p. 9). According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the most representative users were the day laborers and workers classified as “miscellaneous,” accounting for $63.4$% of all users. By contrast, those for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and $15.6$%, respectively (Tokyo Institute for Municipal Research 1926, p. 25). The inexpensive articles pawned suggest that the pawnshops were only used to dealing with very short-run loans for necessities. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which accounts for $12$% of the monthly disposable income of the THBS households (Panel A of Table 3). Thus, the pawnshop was a small lending institution for low-income working-class households who needed credit. Indeed, lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the main articles pawned were inexpensive clothes (Shibuya et al. 1982). Additionally, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates. [^24] The average redemption rate was approximately $94$% in 1920 among $88$ pawnshops in Kyobashi ($401,222$ cases/$427,265$ cases) (Tokyo City Office 1922b, pp. 888–889).
@@
Other lending institutions were money lenders (kinsen kashitsuke gyō) and credit unions (shinyō kumiai). Generally, moneylenders were used by business owners with very high interest rates (Shibuya 2000, pp. 184; 248). Similarly, the coverage of the credit unions was substantially low, possibly a few percentage points of the male workers in the manufacturing sector (Online Appendix A.2).
Informal gifts might have been another type of risk-mitigation institution, although systematic statistics are not available. A household survey on the $185$ factory-worker households in Tokyo suggests that the share of average monthly gifts in total income was approximately $3$% in November 1922 (Social Affairs Division 1925, pp. 58–59). While this means that the gifts comprised a few percentage points of the monthly income, it is unclear whether these informal income transfers functioned as mutual aid.
@@
Finally, life insurance was not designed to compensate for temporary income losses. Generally, life insurance companies provided death insurance, meaning that temporary income reductions were not covered. Moreover, they did not target the working class as customers (Usami 1984, pp. 87–88). Although postal life insurance (kani seimei hoken) operated by the government started in 1916, it only provided life and endowment insurance. Thus, it never covered illness or unemployment (Usami 1984, pp. 102–107).  [^25] There were also several employment agencies that introduced jobs to low-income job-seekers in Tokyo. However, many of them were unscrupulous, and not commonly used at that time (Machida 2016).
@@
To summarize, while there were several lending institutions, few were available to the factory workers. Postal savings might have been the best available device for them. Credit purchase was the most plausible device regarding lending institutions. While pawnshops were accessible, they might not have provided sufficient money as the articles pawned were usually inexpensive. The factory workers in Tsukishima might have combined these devices to manage idiosyncratic income shocks.